A sample of chlorine gas occupies a volume of 946 mL at a

pressure of 726 mmHg. What is the pressure of the gas (in

mmHg) if the volume is reduced at constant temperature to

154 mL?


1
Expert's answer
2022-01-25T11:20:01-0500

P×V = P1×V1 (T=const)

P1 = (P×V)/V1

P1 = (726×0.946) / 0.154 = 594578028 Pa = 4459.714 mmHg

Answer

P1 = 4459.714 mmHg